OpenStudy (anonymous):

I am having trouble figuring out APR using Excel for this one. Anyone know what to do? A used car with a price of $7,880 can be purchased for no money down by making 60 monthly payments of $185.53. Find (a) the total monthly payments, (b) the total finance charge, and (c) the APR.

OpenStudy (phi):

solve for "i" in 7880(1+i/12)^60 = 60*185.53

OpenStudy (phi):

you should get \[ \left( 1+ \frac{i}{12}\right)^{60} = \frac{60 \cdot 185.53}{7880} \] take the 1/60 root of both sides (raise both sides to the 1/60 power) \[ 1+ \frac{i}{12}= \left( \frac{60 \cdot 185.53}{7880}\right)^{\frac{1}{60}}\] can you finish ?